Miller Brothers Hardware paid an annual dividend of $1.75 per share last month. Today, the company announced that future dividends will be increasing by 3.60 percent annually. If you require a 8.1 percent rate of return, how much are you willing to pay to purchase one share of this stock today? 


O $38.54 

O $50.36 

O $48.61 

O $40.29 

O $42.04
Answer #1

Cost of equity = ke = D1 = 8.10% 3.60% $1.81 $40.29 1.81/(8.10% - 3.60%) 1.75*1.036 Price = D1/(ke-g)
Answer #2

ANSWER :


Given :


D0 = 1.75 ($)

g = 3.6 % =  0.036

r = 8.1 % = 0.081


Now, as per DDM at constant growth :


Current value off the stock

= D0 (1 + g) / (r - g)

= 1.75 * (1 + 0.036)/ (0.081 - 0.036)

=  40.29 ($)


Maximum that can be paid per share = 40.29 ($) (ANSWER)



answered by: Tulsiram Garg